OpenStudy (anonymous):

At 1 atm , if in the mercury coloum , mercury was replaced by water the height of the coloum would be?

mathslover (mathslover):

what do u want to calculate @Yahoo!

mathslover (mathslover):

and may i know that what you tried?

OpenStudy (anonymous):

we need to find the height of mercuty colum

mathslover (mathslover):

ok so what did u try ?

OpenStudy (anonymous):

i think there is some formula

OpenStudy (anonymous):

but i cant remember that

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

wait

OpenStudy (anonymous):

pressure = density x g x height

OpenStudy (anonymous):

density of water is 1.00g/cm3

mathslover (mathslover):

pressure is also given 1atm ... g is 9.8 m\s^2

mathslover (mathslover):

solve for h now

OpenStudy (anonymous):

@mathslover but the asnwer should be 10336mm

mathslover (mathslover):

1 meter = how much mili meter ?

OpenStudy (anonymous):

1000

mathslover (mathslover):

1.00g/cm3 = 1g / 1000 mm^3

OpenStudy (anonymous):

@mathslover r u stuck?

mathslover (mathslover):

yes too much :(

mathslover (mathslover):

h1d1g = h2d2g why are we not comparing that of water to mercury ?

OpenStudy (maheshmeghwal9):

sorry dude me no chem:/

mathslover (mathslover):

so u have got the answer

OpenStudy (anonymous):

@mathslover we dont have h1 and h2 sooo

mathslover (mathslover):

ok ...

OpenStudy (anonymous):

Nop in that they have given the height

OpenStudy (anonymous):

i linked to get idea for u

mathslover (mathslover):

wait i am searching for a better idea

OpenStudy (anonymous):

ok

mathslover (mathslover):

1 atmosphere = 760 mmHg

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathslover u got it

mathslover (mathslover):

yes ... some what

OpenStudy (anonymous):

yup can u show that!!!

mathslover (mathslover):

that is all written in comments lolz :( sorry i am trying a lot

OpenStudy (anonymous):

wat u mean?

mathslover (mathslover):

@quarkine is there .. no worries now

OpenStudy (anonymous):

P2-P1= d*g*h  so h can be calculated as we know d(density),g and Pa=P2...so what is the question?? @mathslover :will try..